### v1

@@ 1 Introduction
Quantum enhancement of machine learning attracts much attentions due to provable advantages of certain quantum machine learning (QML) methods over classical counterparts wiebe2012quantum ; aimeur2013quantum ; rebentrost2014quantum ; lloyd2014quantum ; biamonte2017quantum . Of particular interest are QML approaches for data analysis. A family of fruitful results have theoretically shown that tailored quantum models can efficiently learn patterns of synthetic datasets that are hard for classical ones to find liu2021rigorous ; jager2023universal ; Muser2023tos . Therefore, the next step is to quest whether such advantages can be translated into practical machine learning tasks.
@@
A primary candidate of supervised QML models in the near-term is the so-called quantum neural networks schuld2014quest . The key idea of this approach is to utilize the Hilbert space as the feature space for machine learning tasks by exploiting the power of parameterized quantum circuits (PQCs) and data-embedding circuits havlivcek2019supervised ; schuld2019quantum . Due to the mathematical equivalence, learning models based on quantum kernels can also be in the same class of models as the quantum neural networks schuld2021supervised ; quantum neural networks and quantum kernel methods are categorized as explicit and implicit models within the same concept of QML, respectively jerbi2023quantum ; yano2021efficient ; nghiem2021unified ; herman2023expressivity ; sweke2023potential . Focus of this paper is on these supervised QML models.
@@
Thus far, a number of attempts have been made to understand these quantum models analytically. In contrast to the original expectation, some literatures have elucidated that the high expressivity of the models causes trainability problems such as barren plateaus mcclean2018barren ; cerezo2021cost ; holmes2021barren ; holmes2022connecting ; leone2022practical and the exponential concentration huang2021power ; kubler2021inductive ; thanasilp2022exponential ; suzuki2022quantum . This immediately suggests the need to carefully construct quantum models for real-world applications. A remedy for the issue in classical machine learning is feature selection guyon2003introduction ; saeys2007review ; chandrashekar2014survey ; pudjihartono2022review . With the technique, redundant features can be removed and hence the predictive performance improves. Moreover, the model’s interpretability can be enhanced thanks to the reduced features.
@@
We note that feature selection techniques have been studied in the quantum computing community. For example, Quadratic Unconstrained Binary Optimization problems corresponding to the classical feature selection have been introduced, so that quantum computing can deal with it to boost the performance zoufal2023variational ; mucke2023feature ; hellstern2023quantum . Other works use quantum support vector machines with multi-objective genetic algorithms wang2023novel , quantum circuit evolution algorithms albino2023evolutionary and graph-theoretic approach using quantum computers chakraborty2020hybrid for the performance improvement. However, the usability of classical feature selection could be limited for QML tasks. While classical feature selection methods still play a critical role in embedding data on quantum circuits, it does not help understand why some data embedding quantum circuits work better than others. More importantly, when dealing with quantum data tasks such as quantum phase recognition, classical input features do not explicitly appear in input quantum data and thus one cannot employ the classical techniques straightforwardly.
@@
In this work, we propose a feature selection method with a specific focus on QML problems. More specifically, we treat the light-cones structures (i.e., subspaces) of local quantum kernels as quantum features and then select a subset of beneficial ones through training of parameters in the kernels. Note that, by local quantum kernels, we mean a family of quantum kernels that measure local similarity between a pair of data; examples are projected quantum kernels (PQKs) huang2021power and quantum Fisher kernels (QFKs) suzuki2022quantum . Thanks to the extension, the proposed framework is applicable not only to the feature selection of classical inputs, but also to the quantum data tasks that conventional feature selection methods cannot handle. To the best of our knowledge, this is the first work to introduce a QML-oriented feature selection method. Numerical simulations using toy tasks demonstrate the effectiveness of our scheme for four applications: (1) classical feature selection, (2) circuit architecture search for data embedding, (3) the compression of QML models and (4) relevant subspace selection for quantum data tasks. These results indicate that our proposal will encourage practitioners to use QML models for real-world applications. In addition, as elaborated in Conclusion & Discussion section, this scheme is linked to the argument about classical simulability and trainability cerezo2023does . Therefore, our scheme could work as a practical test to see if quantumness is really needed for the QML tasks at hand, while this is left for future work.
@@ 2.1 Framework of Quantum Machine Learning
In case of classical data, the dataset is composed of pairs of classical input $\bm{x}_{i}$ and its target value or label $y_{i}$; an example of the input is a vectorized $d$-dimensional image $\bm{x}\in\mathbb{R}^{d}$. For QML models to deal with the classical data, we have to encode the input to quantum states via embedding quantum circuits. Namely, we apply an embedding circuit $V(\bm{x}_{i})$ to an initial state $\rho_{0}$ to construct data-dependent quantum states $\rho_{i}=V(\bm{x}_{i})\rho_{0}V^{\dagger}(\bm{x}_{i})$. We can also utilize parameterized quantum circuits (PQCs) $U(\bm{\theta})$ with tunable parameters $\bm{\theta}$ to find better features in the space. Typically, we employ data re-uploading technique perez2020data to create quantum states from classical inputs;
@@
In the numerical experiments in Section 3, we use the re-uploading quantum circuits with a brick-like structure cerezo2021cost ; an example for $n=6$ and $L=3$ is shown in Figure 1.
Figure 1: An example of the embedding quantum circuit with $L=3$ for six features. The colored gates represent the light-cone of measurement on $l$-th qubit with $l=3,4$ to construct the PQKs in Eq. (7).
As for quantum data, the dataset constitutes quantum states and their targets, i.e., $\{\rho_{i},y_{i}\}$ schatzki2021entangled ; perrier2022qdataset . In this case, we employ PQCs $U(\bm{\theta})$ to the quantum inputs to find better representation for the tasks;
@@
Previous work defined two approaches of constructing quantum models jerbi2023quantum : explicit models where model predictions are given by measuring a certain Hermitian operator $O$ on quantum states, and implicit models where the inner products of quantum states called quantum kernels are used for predictions. Then, with a training dataset $S=\{\bm{x}_{i},y_{i}\}_{i=1,\ldots,N}$ ($S=\{\rho_{i},y_{i}\}_{i=1,\ldots,N}$) of size $N$, we perform optimization to obtain performant quantum models. For explicit models, we optimize $\bm{\theta}$ to minimize a loss function $\mathcal{L}(\bm{\theta})$ represented as
@@
where $f_{i}(\bm{\theta})=\mathrm{Tr}[\rho_{i}O]$ is the output of explicit model and $l(f_{i}(\bm{\theta}),y_{i})$ is the loss such as squared loss, i.e., $l(f_{i}(\bm{\theta}),y_{i})=(f_{i}(\bm{\theta})-y_{i})^{2}$. On the other hand, the representer theorem kimeldorf1971some ; smola1998learning ; hofmann2008kernel guarantees that an optimal solution of implicit models can be in the form of $f_{i}^{opt}(\bm{x})=\sum_{i}\alpha_{i}^{opt}k(\bm{x}_{i},\bm{x})$ with a kernel function $k(\bm{x},\bm{x^{\prime}})$. A common choice of the kernel for QML is the fidelity quantum kernel havlivcek2019supervised expressed as
@@
Note that the kernel function can include the parameters $\bm{\theta}$ as shown in Eq. (5); the parameters $\bm{\theta}$ can also be optimized separately JMLR:v13:cortes12a or simultaneously Sydorov_2014_CVPR . Basically, classical optimizers are employed for both models in the optimization.
@@ 2.2 Light-Cone Feature Selection
We here extend the idea of classical feature selection guyon2003introduction to propose its counterpart for QML. The classical feature selection methods can be categorized into three types: the filter, wrapper, and embedded methods. The filter approach performs selections based on the relationship between features and the targets. The wrapper approach evaluates the performance of the model for different subsets of features and then select features based on a performance metric. Lastly, the embedded approach provides important features via training of models such as LASSO. See, e.g., Ref. saeys2007review ; chandrashekar2014survey ; pudjihartono2022review for more comprehensive review of classical feature selections.
Our proposal is similar in spirit to the embedded approach; we select relevant local light-cone of quantum models through training of local quantum kernels. By local quantum kernels, we mean quantum kernels which measure local similarity between a pair of data. Note that fidelity quantum kernels in Eq. (5) measure global similarity and thus do not fall into the category. The examples are projected quantum kernels (PQKs) huang2021power and simplified quantum Fisher kernels (QFKs) suzuki2022quantum defined as
@@
These local quantum kernels can be expressed as summation of local terms, $k(\bm{x},\bm{x^{\prime}})=\sum_{l}\lambda_{l}k^{(l)}(\bm{x},\bm{x^{\prime}})$, where each local kernel $k^{(l)}(\bm{x},\bm{x^{\prime}})$ has the corresponding light-cone, i.e, the subspace of quantum circuits (or quantum states) that are relevant to a certain operation like a local gate operation and measurement. For example, as illustrated in Figure 1, the light-cone obtained by measurement on $l$-th qubit ($l=3,4$) for the PQK is denoted by the colored gates. We regard the light-cones induced by local quantum kernels as features and then select important ones through training of parameters $(\bm{\theta},\bm{\lambda})$ in quantum kernels.
As for the training of quantum kernels, we maximize the kernel target alignment (KTA) cristianini2001kernel defined as
@@
We note that the KTA is a common choice as an objective function for training kernels in both classical JMLR:v13:cortes12a ; cristianini2001kernel and quantum cases kubler2021inductive ; thanasilp2022exponential ; miyabe2023quantum . In this work, the optimization of $(\bm{\theta},\bm{\lambda})$ is done in two steps: (1) with $\bm{\lambda}$ fixed, $\bm{\theta}$ are optimized via the Adam optimizer by computing the gradients of Eq. (9) with respect to $\bm{\theta}$, and then, (2) $\bm{\lambda}$ are optimized following the approach proposed in Ref. JMLR:v13:cortes12a :
@@
This two-step process is iteratively carried out until the value of the KTA stops changing significantly. The optimization can be done efficiently, as the first step relies on gradient-based algorithms and the solution of the second step can be obtained by solving a quadratic program. In our numerical experiments, Eq. (11) in the second step is solved using CVXOPT library andersen2011 ; cvxopt . After the optimization, we select relevant light-cones based on the amplitude of parameters $\bm{\lambda}$; that is, we interpret the local kernel $k^{(l)}(\bm{x},\bm{x^{\prime}})$ whose parameter $\lambda_{l}$ has a large absolute value as important.
We remark that our proposal is more versatile than classical feature selection methods in QML tasks. This scheme can be used not only for feature selection of classical features as the conventional methods do, but also for seeking relevant subspace of embedding quantum circuits. More importantly, the proposed method is applicable to quantum data tasks where classical features do not appear and hence conventional ones cannot work. In addition, while some previous works proposed methods for feature selection using quantum computing devices, all of them focus on improvement of the classical feature selection zoufal2023variational ; mucke2023feature ; hellstern2023quantum ; wang2023novel ; chakraborty2020hybrid ; albino2023evolutionary . Therefore, to the best of our knowledge, this work is the first to propose a QML-oriented feature selection scheme.
@@ 3 Numerical Demonstration
In what follows, we perform numerical simulations to demonstrate efficacy and versatility of our proposal. More specifically, we deal with four application tasks: (1) feature selection of classical inputs, (2) circuit architecture search for data embedding, (3) compression of QML models and (4) subspace feature selection for quantum data. For all the simulations shown below, PennyLane library bergholm2022pennylane is used for quantum circuit simulation. Also, in the learning tasks, we used 80 data points for both training and test, where models’ performance are checked for different 25 settings of data samples and initial parameters $\bm{\theta}$; five sets of data samples and five different parameters are prepared and then we tried all combinations of these settings. As for local quantum kernels, we set a uniform weights to the initial parameters $\bm{\lambda}$, i.e., $\lambda_{l}=1/$(the number of terms in the local quantum kernel) for all $l$. The learning rate and the number of iterations in the Adam optimization step are set to 0.01 and 500, respectively.
@@ 3.1 Feature Selection of Classical Inputs
where $w_{\lambda_{l}}(\mu)$ is the frequency for the classical feature $\mu$ to appear in the $l$-th local quantum kernel $k^{(l)}(\bm{x},\bm{x’})$, and $\mathcal{N}=\sum_{\mu}\tilde{P}(\mu)$ with $\tilde{P}(\mu)=\sum_{l}w_{\lambda_{l}}(\mu)\lambda_{l}$ is the normalization constant. Recall that $\bm{\lambda}$ satisfy $\sum_{l}\lambda_{l}=1$ with $\lambda_{l}\geq 0$. In case of the light-cone structure shown in Figure 1 (i.e., $\lambda_{l}$ with $l=3$ or $4$ for PQKs), $w_{\lambda_{l}}(\mu)=1$ for $\mu=x_{1},x_{6}$, $w_{\lambda_{l}}(\mu)=2$ for $\mu=x_{2},x_{5}$, and $w_{\lambda_{l}}(\mu)=3$ for $\mu=x_{3},x_{4}$. We notice that one can take another quantity as the importance score, so long as it tells the relevance of features in the tasks.
As a toy example, we consider a so-called parity-based binary classification task introduced in Ref. jerbi2023quantum The goal of this task is to classify $d$-dimensional binary input $\bm{x}\in\{-1,1\}^{d}$ whose label is determined based on its parity for a certain subset $A_{p}\subset\{1,\ldots,d\}$; namely, the input’s label is assigned according to $y(\bm{x})=\prod_{i\in A_{p}}x_{i}$. We chose this task because important classical features are clearly $A_{p}$ and hence we can interpret the performance of our proposal easily. In the following, we consider a dataset with $A_{p}=\{1,2\}$, which we denote as parity(1,2). Here, the number of qubits $n$ is equal to the dimension of inputs $d$ and we set $d=8$. Also, the data re-uploading quantum circuits with the depth $L=3$ is used.
@@
Figure 2 shows the importance scores of the local quantum kernels for the parity(1,2) dataset. Features whose importance scores are higher than the threshold $P_{th}=0.1$ for QFKs are $x_{1}$ and $x_{2}$, suggesting our scheme with QFKs can pinpoint key features for the task. As for PQKs, features $x_{1}$, $x_{2}$, $x_{3}$ and $x_{8}$ witness the scores larger than $P_{th}$. Hence, the PQKs can also select relevant features on the dataset. The features chosen by the PQKs include more features than the minimum required. This is because the number of classical features in the light-cone induced by a local measurement gets larger as the depth increases suzuki2023effect . We note that the light-cone of $l$-th term of QFKs is determined by the position of gates with the local parameter $\theta_{l}$ suzuki2022quantum , and thus only necessary features can be selected in this case. Actually, the performance difference between these two kernels can be seen by looking at the KTAs; the values of the KTAs for PQKs and QFKs are 0.605 and 0.987, respectively. Recall that the KTA of the best kernel outputs one, and the larger value of the quantity indicates the better performance. Hence, the performance gap between the PQK and the QFK would be enlarged as the depth increases. Yet, we could resolve the problem in the PQKs, e.g., by using different feature scores.

### v2

@@ 1 Introduction
Figure 1: Summary of this work. Our proposed scheme selects relevant light-cone features for QML tasks through the training of parameters $(\bm{\theta},\bm{\lambda})$ in local quantum kernels (note that the local term $k^{(l)}(\bm{x},\bm{x^{\prime}})$ includes tunable parameters $\bm{\theta}$). With the light-cone feature selection, we can handle not only the feature selection for classical data, but also the subspace selection for quantum data tasks, which conventional feature selection techniques cannot deal with.
@@
Quantum enhancement of machine learning attracts much attentions due to provable advantages of certain quantum machine learning (QML) methods over classical counterparts [1, 2, 3, 4, 5]. Of particular interest are QML approaches for data analysis. A family of fruitful results have theoretically shown that tailored quantum models can efficiently learn patterns of synthetic datasets that are hard for classical ones to find [6, 7, 8]. Therefore, the next step is to quest whether such advantages can be translated into practical machine learning tasks.
@@
A primary candidate of supervised QML models in the near-term is the so-called quantum neural networks [9]. The key idea of this approach is to utilize the Hilbert space as the feature space for machine learning tasks by exploiting the power of parameterized quantum circuits (PQCs) and data-embedding circuits [10, 11]. Due to the mathematical equivalence, learning models based on quantum kernels can also be in the same class of models as the quantum neural networks [12]; quantum neural networks and quantum kernel methods are categorized as explicit and implicit models within the same concept of QML, respectively [13, 14, 15, 16, 17]. Focus of this paper is on these supervised QML models.
@@
Thus far, a number of attempts have been made to understand these quantum models analytically. In contrast to the original expectation, some literatures have elucidated that the high expressivity of the models causes trainability problems such as barren plateaus [18, 19, 20, 21, 22] and the exponential concentration [23, 24, 25, 26]. This immediately suggests the need to carefully construct quantum models for real-world applications. A remedy for the issue in classical machine learning is feature selection [27, 28, 29, 30]. With the technique, redundant features can be removed and hence the predictive performance improves. Moreover, the model’s interpretability can be enhanced thanks to the reduced features.
@@
We note that feature selection techniques have been studied in the quantum computing community. For example, Quadratic Unconstrained Binary Optimization problems corresponding to the classical feature selection have been introduced, so that quantum computing can deal with it to boost the performance [31, 32, 33]. Other works use quantum support vector machines with multi-objective genetic algorithms [34], quantum circuit evolution algorithms [35] and graph-theoretic approach using quantum computers [36] for the performance improvement. However, the usability of classical feature selection could be limited for QML tasks. While classical feature selection methods still play a critical role in embedding data on quantum circuits, it does not help understand why some data embedding quantum circuits work better than others. More importantly, when dealing with quantum data tasks such as quantum phase recognition, classical input features do not explicitly appear in input quantum data and thus one cannot employ the classical techniques straightforwardly.
@@
In this work, we propose a feature selection method with a specific focus on QML problems. More specifically, we treat the light-cone structures (i.e., subspaces) of local quantum kernels as quantum features and then select a subset of beneficial ones through training of parameters in the kernels. Note that, by local quantum kernels, we mean a family of quantum kernels that measure local similarity between a pair of data; examples are projected quantum kernels (PQKs) [23] and quantum Fisher kernels (QFKs) [26]. Thanks to the extension, the proposed framework is applicable not only to the feature selection of classical inputs, but also to the quantum data tasks that conventional feature selection methods cannot handle. To the best of our knowledge, this is the first work to introduce a QML-oriented feature selection method. Numerical simulations using toy tasks demonstrate the effectiveness of our scheme for four applications: (1) classical feature selection, (2) circuit architecture search for data embedding, (3) the compression of QML models and (4) relevant subspace selection for quantum data tasks. These results indicate that our proposal will encourage practitioners to use QML models for real-world applications. Figure 1 summarizes our work. In addition, as elaborated in Conclusion & Discussion section, this scheme is linked to the argument about classical simulability and trainability [37]. Therefore, our scheme could work as a practical test to see if quantumness is really needed for the QML tasks at hand, while this is left for future work.
@@ 2.1 Framework of Quantum Machine Learning
In case of classical data, the dataset is composed of pairs of classical input $\bm{x}_{i}$ and its target value or label $y_{i}$; an example of the input is a vectorized $d$-dimensional image $\bm{x}\in\mathbb{R}^{d}$. For QML models to deal with the classical data, we have to encode the input to quantum states via embedding quantum circuits. Namely, we apply an embedding circuit $V(\bm{x}_{i})$ to an initial state $\rho_{0}$ to construct data-dependent quantum states $\rho_{i}=V(\bm{x}_{i})\rho_{0}V^{\dagger}(\bm{x}_{i})$. We can also utilize parameterized quantum circuits (PQCs) $U(\bm{\theta})$ with tunable parameters $\bm{\theta}$ to find better features in the space. Typically, we employ data re-uploading technique [38] to create quantum states from classical inputs;
@@
In the numerical experiments in Section 3, we use the re-uploading quantum circuits with a brick-like structure [19]; an example for $n=6$ and $L=3$ is shown in Figure 2.
Figure 2: Examples of embedding quantum circuits with the depth $L=3$ and six features $\bm{x}=[x_{1},\ldots,x_{6}]$ for the PQK and the QFK. The colored gates represent the light-cone of these quantum kernels; each shows the light-cone of measurement on $l$-th qubit with $l=3$ to construct the PQKs in Eq. (7), and that of the gate including the $l$-th parameter $l=\theta\in\bm{\theta}_{2,1}$ for the QFK in Eq. (8).
As for quantum data, the dataset consists of quantum states and their targets, i.e., $\{\rho_{i},y_{i}\}$ [39, 40]. In this case, we employ PQCs $U(\bm{\theta})$ to the quantum inputs to find better representation for the tasks;
@@
Previous work defined two approaches of constructing quantum models [13]: explicit models where model predictions are given by measuring a certain Hermitian operator $O$ on quantum states, and implicit models where the inner products of quantum states called quantum kernels are used for predictions. Then, with a training dataset $S=\{\bm{x}_{i},y_{i}\}_{i=1,\ldots,N}$ ($S=\{\rho_{i},y_{i}\}_{i=1,\ldots,N}$) of size $N$, we perform optimization to obtain performant quantum models. For explicit models, we optimize $\bm{\theta}$ to minimize a loss function $\mathcal{L}(\bm{\theta})$ represented as
@@
where $f_{i}(\bm{\theta})=\mathrm{Tr}[\rho_{i}O]$ is the output of explicit model and $l(f_{i}(\bm{\theta}),y_{i})$ is the loss such as squared loss, i.e., $l(f_{i}(\bm{\theta}),y_{i})=(f_{i}(\bm{\theta})-y_{i})^{2}$. On the other hand, the representer theorem [41, 42, 43] guarantees that an optimal solution of implicit models can be in the form of $f_{i}^{opt}(\bm{x})=\sum_{i}\alpha_{i}^{opt}k(\bm{x}_{i},\bm{x})$ with a kernel function $k(\bm{x},\bm{x^{\prime}})$. A common choice of the kernel for QML is the fidelity quantum kernel [10] expressed as
@@
Note that the kernel function can include the parameters $\bm{\theta}$ as shown in Eq. (5); the parameters $\bm{\theta}$ can also be optimized separately [44] or simultaneously [45]. Basically, classical optimizers are employed for both models in the optimization.
@@ 2.2 Light-Cone Feature Selection
We here extend the idea of classical feature selection [27] to propose its counterpart for QML. The classical feature selection methods can be categorized into three types: the filter, wrapper, and embedded methods. The filter approach performs selections based on the relationship between features and the targets. The wrapper approach evaluates the performance of the model for different subsets of features and then select features based on a performance metric. Lastly, the embedded approach provides important features via training of models such as LASSO. See, e.g., Ref. [28, 29, 30] for more comprehensive review of classical feature selections.
Our proposal is similar in spirit to the embedded approach; we select relevant local light-cone of quantum models through training of local quantum kernels. By local quantum kernels, we mean quantum kernels which measure local similarity between a pair of data. Note that fidelity quantum kernels in Eq. (5) measure global similarity and thus do not fall into the category. The examples are projected quantum kernels (PQKs) [23] and simplified quantum Fisher kernels (QFKs) [26] defined as
@@
These local quantum kernels can be expressed as summation of local terms, $k(\bm{x},\bm{x^{\prime}})=\sum_{l}\lambda_{l}k^{(l)}(\bm{x},\bm{x^{\prime}})$, where each local kernel $k^{(l)}(\bm{x},\bm{x^{\prime}})$ has the corresponding light-cone, i.e, the subspace of quantum circuits (or quantum states) that are relevant to a certain operation like a local gate operation and measurement. For example, as illustrated in Figure 2, the light-cones obtained by measurement on $l$-th qubit ($l=3$) for the PQK and a gate operation including $l$-th parameter ($l=\theta\in\bm{\theta}_{2,1}$) for the QFK are represented by the colored gates. We regard the light-cones induced by local quantum kernels as features and then select important ones through training of parameters $(\bm{\theta},\bm{\lambda})$ in quantum kernels.
As for the training of quantum kernels, we maximize the kernel target alignment (KTA) [46] defined as
@@
We note that the KTA is a common choice as an objective function for training kernels in both classical [44, 46] and quantum cases [24, 25, 47]. In this work, the optimization of $(\bm{\theta},\bm{\lambda})$ is done in two steps: (1) with $\bm{\lambda}$ fixed, $\bm{\theta}$ are optimized via the Adam optimizer by computing the gradients of Eq. (9) with respect to $\bm{\theta}$, and then, (2) $\bm{\lambda}$ are optimized following the approach proposed in Ref. [44]:
@@
This two-step process is iteratively carried out until the value of the KTA stops changing significantly. The optimization can be done efficiently, as the first step relies on gradient-based algorithms and the solution of the second step can be obtained by solving a quadratic program. In our numerical experiments, Eq. (11) in the second step is solved using CVXOPT library [48, 49]. After the optimization, we select relevant light-cones based on the amplitude of parameters $\bm{\lambda}$; that is, we interpret the local kernel $k^{(l)}(\bm{x},\bm{x^{\prime}})$ whose parameter $\lambda_{l}$ has a large absolute value as important.
We remark that our proposal is more versatile than classical feature selection methods in QML tasks. This scheme can be used not only for feature selection of classical features as the conventional methods do, but also for seeking relevant subspace of embedding quantum circuits. More importantly, the proposed method is applicable to quantum data tasks where classical features do not appear and hence conventional ones cannot work. In addition, while some previous works proposed methods for feature selection using quantum computing devices, all of them focus on improvement of the classical feature selection [31, 32, 33, 34, 36, 35]. Therefore, to the best of our knowledge, this work is the first to propose a QML-oriented feature selection scheme.
@@ 3 Numerical Demonstration
In what follows, we perform numerical simulations to demonstrate efficacy and versatility of our proposal. More specifically, we deal with four application tasks: (1) feature selection of classical inputs, (2) circuit architecture search for data embedding, (3) compression of QML models and (4) subspace feature selection for quantum data. For all the simulations shown below, PennyLane library [50] is used for quantum circuit simulation. Also, in the learning tasks, we used 80 data points for both training and test, where models’ performance are checked for different 25 settings of data samples and initial parameters $\bm{\theta}$; five sets of data samples and five different parameters are prepared and then we tried all combinations of these settings. As for local quantum kernels, we set a uniform weights to the initial parameters $\bm{\lambda}$, i.e., $\lambda_{l}=1/$(the number of terms in the local quantum kernel) for all $l$. The learning rate and the number of iterations in the Adam optimization step are set to 0.01 and 500, respectively.
@@ 3.1 Feature Selection of Classical Inputs
where $w_{\lambda_{l}}(\mu)$ is the frequency for the classical feature $\mu$ to appear in the $l$-th local quantum kernel $k^{(l)}(\bm{x},\bm{x\textquoteright})$, and $\mathcal{N}=\sum_{\mu}\tilde{P}(\mu)$ with $\tilde{P}(\mu)=\sum_{l}w_{\lambda_{l}}(\mu)\lambda_{l}$ is the normalization constant. Recall that $\bm{\lambda}$ satisfy $\sum_{l}\lambda_{l}=1$ with $\lambda_{l}\geq 0$. In case of the light-cone structure of the PQK shown in Figure 2 (i.e., $\lambda_{l}$ with $l=3$ depicted in the upper quantum circuit), $w_{\lambda_{l}}(\mu)=1$ for $\mu=x_{1},x_{6}$, $w_{\lambda_{l}}(\mu)=2$ for $\mu=x_{2},x_{5}$, and $w_{\lambda_{l}}(\mu)=3$ for $\mu=x_{3},x_{4}$. We notice that one can take another quantity as the importance score, so long as it tells the relevance of features in the tasks.
As a toy example, we consider a so-called parity-based binary classification task introduced in Ref. [13]. The goal of this task is to classify $d$-dimensional binary input $\bm{x}\in\{-1,1\}^{d}$ whose label is determined based on its parity for a certain subset $A_{p}\subset\{1,\ldots,d\}$; namely, the input’s label is assigned according to $y(\bm{x})=\prod_{i\in A_{p}}x_{i}$. We chose this task because important classical features are clearly $A_{p}$ and hence we can interpret the performance of our proposal easily. In the following, we consider a dataset with $A_{p}=\{1,2\}$, which we denote as parity(1,2). Here, the number of qubits $n$ is equal to the dimension of inputs $d$ and we set $d=8$. Also, the data re-uploading quantum circuits with the depth $L=3$ is used.
@@
Figure 3 shows the importance scores of the local quantum kernels for the parity(1,2) dataset. Features whose importance scores are higher than the threshold $P_{th}=0.1$ for QFKs are $x_{1}$ and $x_{2}$, suggesting our scheme with QFKs can pinpoint key features for the task. As for PQKs, features $x_{1}$, $x_{2}$, $x_{3}$ and $x_{8}$ witness the scores larger than $P_{th}$. Hence, the PQKs can also select relevant features on the dataset. The features chosen by the PQKs include more features than the minimum required. This is because the number of classical features in the light-cone induced by a local measurement gets larger as the depth increases [51]. We note that the light-cone of $l$-th term of QFKs is determined by the position of gates with the local parameter $\theta_{l}$ [26], and thus only necessary features can be selected in this case. Actually, the performance difference between these two kernels can be seen by looking at the KTAs; the values of the KTAs for PQKs and QFKs are 0.605 and 0.987, respectively. Recall that the KTA of the best kernel outputs one, and the larger value of the quantity indicates the better performance. Hence, the performance gap between the PQK and the QFK would be enlarged as the depth increases. Yet, we could resolve the problem in the PQKs, e.g., by using different feature scores.